Solve for x
4.2^x = (0.5)^x-2
and
8^-x = 2.4^x-1

Answer by Alan3354(69443)

Solve for x
4.2^x = (0.5)^x-2
x*log(4.2) = (x-2)*log(0.5) = x*log(0.5) - 2log(0.5)
x*log(4.2) - x*log(0.5) = -2log(0.5)
x = -2log(0.5)/(log(4.2 - log(0.5))
Calculator work
-------------------
and
8^-x = 2.4^x-1
Do it like the 1st one.